# Looking through the Kerr disk

### v1

@@ Appendix G Gravitational redshift
To create images of what an actual physical observer might see, one task is to consider the redshift along the geodesic. In this section, we discuss the gravitational redshift and comment on the Doppler shift due to the relative motion of the source and observer, which we want to consider. The calculation is done using Boyer-Lindquist coordinates $(t_{BL},r_{BL},\theta_{BL},\varphi_{BL})$ of the Kerr spacetime [visser_kerr_2008]. For brevity, we will denote them by $(t,r,\theta,\varphi)$ in this section.
@@
Note that this does not necessarily correspond to an observer with zero angular momentum at any radial position $r_{o}$, who, due to the dragging of frames by the rotation of the black hole, does not have a constant azimuthal coordinate but rather rotates with an angular velocity [bardeen_timelike_1972]
@@
We calculate the gravitational redshift in the same manner as one does the calculation in the Schwarzschild case (see, e.g., [chrusciel_elements_2020]). The frequency of the emitted light, as measured by an observer with four-velocity $U_{a}$ ($a\in\{o,s\}$, where $o$ denotes the considered observer and $s$ denotes the light source), is given by
@@
$g(\partial_{t},K)=g_{tt}\,\kappa\,\dot{t}+g_{t\varphi}\,\kappa\,\dot{\varphi}=\operatorname{const}\,$ (126)

### v2

@@ Appendix G Gravitational redshift
To create images of what an actual physical observer might see, one task is to consider the redshift along the geodesic. In this section, we discuss the gravitational redshift and comment on the Doppler shift due to the relative motion of the source and observer, which we want to consider. The calculation is done using Boyer-Lindquist coordinates $(t_{BL},r_{BL},\theta_{BL},\varphi_{BL})$ of the Kerr spacetime [24]. For brevity, we will denote them by $(t,r,\theta,\varphi)$ in this section.
@@
Note that this does not necessarily correspond to an observer with zero angular momentum at any radial position $r_{o}$, who, due to the dragging of frames by the rotation of the black hole, does not have a constant azimuthal coordinate but rather rotates with an angular velocity [3]
@@
We calculate the gravitational redshift in the same manner as one does the calculation in the Schwarzschild case (see, e.g., [6]). The frequency of the emitted light, as measured by an observer with four-velocity $U_{a}$ ($a\in\{o,s\}$, where $o$ denotes the considered observer and $s$ denotes the light source), is given by
@@
[…]